[Hallc_running] RC report 04/09/2019

gabriel at jlab.org gabriel at jlab.org
Tue Apr 9 15:14:51 EDT 2019

Dear all-
This is the RC report for 04/09/2019:

- no RC meeting today
- beam studies all day yesterday. beam restored ~16:20
- weather-related instabilities during the evening as thunderstorms rolled in
- switched to the Al dummy and completed all running on that target
- returned to LH2 and ran throughout the owl and day shifts. Beam has been
a little trippy, with MCC occasionally lowering the current to 60 uA in an
attempt to reduce the number of trips/hour.
- up 15:00 the total number of K-Lambda events at the current setting was
- assuming we will hit the lower limit for the "want" events at this
setting during the swing/owl shift then the next kinematic change should
occur sometimes tomorrow, 04/10/2019, morning.

- at 15:30 today Bill Lee will take over as RC.

It's been a pleasure serving as RC this past (almost a) week.
Gabriel Niculescu

More information about the Hallc_running mailing list